OpenStudy (anonymous):

does anyone know how to write y=-2x^2-4x+3 in intercept form using the quadratic formula? :)

OpenStudy (shamil98):

you need to find the x intercepts since the formula for intercept form is y= a(x-p)(x-k) where x = p x =k are the x-intercepts for your graph

OpenStudy (anonymous):

okay makes sense, so if i had an equation already in intercept form like y=(x+2)(x-3) how would I write that in standard form?

OpenStudy (anonymous):

oh okay i thought that was vertex form

OpenStudy (shamil98):

oh oops wrong one, i mean standard form is ... f(x) = ax^2 + bx + c

OpenStudy (shamil98):

y = (x+2)(x-3) you just multiply them to get that

OpenStudy (anonymous):

oh okay, yes thats what i have :) i just dont know how to change it into different forms

OpenStudy (shamil98):

wait, my alg 2 is rusty, hold on i need to look this up..

OpenStudy (anonymous):

would it be x^2-x-6?

OpenStudy (shamil98):

ah yes, i was right, yeah that is your standard form

OpenStudy (anonymous):

okay, so how do i write that same equation in vertex form?

OpenStudy (shamil98):

-b/2a = x coordinate for your vertex

OpenStudy (anonymous):

so x would be 1/2?

OpenStudy (shamil98):

Yes, then plug that value in into your original function x^2 - x - 6 = y for the y value of the vertex

OpenStudy (shamil98):

(1/2)^2 - 1/2 - 6 = y

OpenStudy (shamil98):

so 1/4 - 2/4 - 6 =y y = -1/4 - 6 y = -25/4

OpenStudy (anonymous):

so what would i do now to get the vertex form?

OpenStudy (shamil98):

your vertex is (1/2,-25/4) vertex form is y = a(x-h)^2 + k y = (x-1/2)^2 - 25/4

OpenStudy (anonymous):

thank you so much!